1. Sequential (Step) Method of Support Department Cost Allocation

    Valron Company has two support departments, Human Resources and General Factory, and two producing departments, Fabricating and Assembly.

    Support Departments Producing Departments
    Human
    Resources
    General
    Factory
    Fabricating Assembly
    Direct costs $170,000    $330,000    $114,600    $91,000   
    Normal activity:
       Number of employees —    50    70    160   
       Square footage 1,100    —    5,900    13,000   

    Resources Department are allocated on the basis of number of employees, and the costs of General Factory are allocated on the basis of square footage. Now assume that Valron Company uses the sequential method to allocate support department costs. The support departments are ranked in order of highest cost to lowest cost.

    Required:

    1. Calculate the allocation ratios (rounded to four significant digits) for the four departments using the sequential method. If an amount is zero, enter "0". Use the rounded values for subsequent calculations.

    Proportion of Driver Used by
    Human Resources General Factory Fabricating Assembly
    Human Resources
    General Factory

    2. Using the sequential method, allocate the costs of the Human Resources and General Factory departments to the Fabricating and Assembly departments. If an amount is zero, enter"0". Round your answers to the nearest dollar.

    Support Departments Producing Departments
    Human Resources General Factory Fabricating Assembly
    Direct costs $ $ $ $
    Allocate:
      General Factory            
      Human Resources            
    Total after allocation $ $ $ $
Answer #1

Answers

  • Requirement 1

Proportion of Driver Used by

Human Resources

General Factory

Fabricating

Assembly

Human Resources

0.3043 [70/(70+160)]

0.6957 [160/(70+160)]

General Factory

0.0550 [1100/(1100 + 5900 + 13000)]

0.2950 [5900/(1100 + 5900 + 13000)]

0.6500 [5900/(13000 + 5900 + 13000)]

  • Requirement 2

Support Departments

Producing Departments

Human Resources

General Factory

Fabricating

Assembly

Direct costs

$170,000

$330,000

$114,600

$91,000

Allocate:

  General Factory

$18,150

($330,000)

$97,350

$214,500

  Human Resources

($188,150)

$0

$57,254

$130,896

Total after allocation

$0

$0

$0

$269,204

$436,396

--Working

Support Departments

Producing Departments

Human Resources

General Factory

Fabricating

Assembly

Direct costs

170000

330000

114600

91000

Allocate:

  General Factory

=330000*0.055

-330000

=330000*0.295

=330000*0.65

  Human Resources

=-170000-18150

0

=188150*0.3043

=188150*0.6957